if a force of 80N extends a spring of natural lenght 8m by 0.4m what will be the lenght of the spring when the force applied is 100N.
1
Expert's answer
2013-01-10T09:48:31-0500
Hooke's law:
F=k*ΔL
k - is a constant called the rate or spring constant
ΔL - is the displacement of the spring's end from its equilibrium position
F1 = k*ΔL1
F2 = k*ΔL2
F1/F2=ΔL1/ΔL2
ΔL2=ΔL1*F2/F1 = 0.4m *100/80 =0.5 m
total length:
L=L0+ΔL2 = 8 + 0.5 = 8.5 m
L0 - natural length
Answer: L=8.5 m